# Quantitative estimates for singularity for conjugate equations driven by linear fractional transformations

### v3

@@ 2.2. Weak contraction
Hereafter, we assume more additional conditions for the D-systems $\{f_{i}\}_{i}$ and $\{g_{i}\}_{i}$. We say that a function $h:[0,1]\to[0,1]$ is a contraction on $[0,1]$ if
$\textup{Lip}(h):=\sup_{x\neq y}\left|\frac{h(x)-h(y)}{x-y}\right|<1.$
$\textup{Lip}(h)$ is called the Lipschitz constant of $h$. We say that a function $h:[0,1]\to[0,1]$ is a weak contraction on $[0,1]$ if
$\phi_{h}(t):=\lim_{s\to t+0}\sup\left\{|h(x)-h(y)|:|x-y|\leq s\right\}<t,\ \ t>0.$
The following is well-known. See [5, 14].
@@
Let $h$ be a weak contraction. Then, $\lim_{n\to\infty}\phi_{h}^{n}(t)=0$ for every $t>0$ and $|h(x)-h(y)|\leq\phi_{h}(|x-y|)$ for $x,y\in[0,1]$.
@@
Let
$\omega_{h}(s):=\sup\left\{|h(x)-h(y)|:|x-y|\leq s\right\}.$
Since this function is increasing, $\phi_{h}$ is right-continuous and increasing on $[0,1]$. Hence, $|h(x)-h(y)|\leq\lim_{s\to|x-y|+0}\phi_{h}(s)=\phi_{h}(|x-y|)$.
By the assumption, for every $t>0$, $(\phi_{h}^{n}(t))_{n}$ is strictly decreasing. There exists the limit $\phi^{\infty}_{h}(t):=\lim_{n\to\infty}\phi_{h}^{n}(t)$. By the right-continuity of $\phi_{h}$, $\phi_{h}\left(\phi^{\infty}_{h}(t)\right)=\phi^{\infty}_{h}(t)$. Hence, $\phi^{\infty}_{h}(t)=0$. ∎
@@
Let $\{h_{i}\}_{i}$ be a compatible system on $[0,1]$ such that each $h_{i}$ is a weak contraction on $[0,1]$. Then, $\{h_{i}\}_{i\in\mathcal{I}_{N}}$ is a D-system.
We say that a compatible system $\{h_{i}\}_{i}$ on $[0,1]$ is an (weakly) contractive system on $[0,1]$ if each $h_{i}$ is an (weak) contraction on $[0,1]$. Every weakly contractive system is a D-system.
@@
Let $\phi:=\max_{i\in\mathcal{I}_{N}}\phi_{h_{i}}$. Then, $\phi$ is right-continuous [^1] The assumption that $N$ is finite would be important. on $[0,1]$, $\phi(t)<t$ for every $t>0$, and $\max_{i\in\mathcal{I}_{N}}|h_{i}(x)-h_{i}(y)|\leq\phi(|x-y|)$, $x,y\in[0,1]$. Then, we see that $\lim_{n\to\infty}\phi^{n}(t)=0$ for every $t>0$. Hence,
@@
$\lim_{n\to\infty}\max_{i_{1},\cdots,i_{n}\in\mathcal{I}_{N}}|h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1)-h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0)|=0.$
@@
$[0,1]=\bigcup_{i_{1},\cdots,i_{n}\in\mathcal{I}_{N}}\left[h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0),h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1)\right],$
@@ 2.3. Linear fractional transforms
We say that a family of functions $\{h_{i}\}_{i}$ is an LF system on $[0,1]$ if $h_{i}(x):=\dfrac{a_{i}x+b_{i}}{c_{i}x+d_{i}},x\in[0,1]$, for each $i\in\mathcal{I}_{N}$, where $a_{i},b_{i},c_{i},d_{i},i\in\mathcal{I}_{N}$, are real numbers such that
(1) (well-defined) $c_{i}+d_{i}>0$, $d_{i}>0$, $i\in\mathcal{I}_{N}$.
(2) (strictly increasing, weak contraction)
@@
(3) (compatibility) $b_{0}=0$, $\dfrac{a_{i}+b_{i}}{c_{i}+d_{i}}=\dfrac{b_{i+1}}{d_{i+1}}$, $i\in\mathcal{I}_{N}$, where we let $\dfrac{b_{N}}{d_{N}}:=1$.
For condition (2), we may consider a stronger condition that
@@
Since $h_{i}$ is invariant if we multiply a positive constant by each of $a_{i},b_{i},c_{i}$ and $d_{i}$, we may assume that $d_{i}=1$, or, $a_{i}d_{i}-b_{i}c_{i}=1$.
@@
Every LF system is a weakly contractive system. If (2’) holds instead of (2), then, it is a contractive system.
@@
Let $\{h_{i}\}_{i\in\mathcal{I}_{N}}$ be an LF-system. We assume that $d_{i}=1$ for every $i$. By Lemma 2.6, it suffices to show that each $h_{i}$ is a weak contraction. We first remark that
$|h_{i}(x)-h_{i}(y)|=\frac{a_{i}-b_{i}c_{i}}{(c_{i}x+1)(c_{i}y+1)}|x-y|.$
@@
(iii) Assume $c_{i}=0$. Then, $0<a_{i}\leq 1$ and $|h_{i}(x)-h_{i}(y)|=a_{i}|x-y|$. By the assumption for $a_{i},b_{i},c_{i}$, $h_{j}(1)>h_{j}(0)$ for every $j$. Hence, by $N\geq 2$, $a_{i}=h_{i}(1)-h_{i}(0)<1$. ∎
@@
Let $(\mathcal{I}_{N})^{*}:=\cup_{n\in\mathbb{N}}(\mathcal{I}_{N})^{n}$. We denote the length of $\sigma\in(\mathcal{I}_{N})^{*}$ by $|\sigma|$. For an LF system $\{h_{i}\}_{i}$ and $\sigma=(\sigma_{1},\cdots,\sigma_{n})\in(\mathcal{I}_{N})^{*}$, we let
$h_{\sigma}(x):=h_{\sigma_{1}}\circ\cdots\circ h_{\sigma_{n}}(x),\ x\in[0,1].$
@@
Let $\{h_{i}\}_{i}$ be an LF system. Then, there exists $c\in(0,1)$ such that for every $\sigma\in(\mathcal{I}_{N})^{*}$,
@@
Let $A_{i}:=\begin{pmatrix}a_{i}&b_{i}\\ c_{i}&d_{i}\end{pmatrix}$, $i\in\mathcal{I}_{N}$. We assume that $a_{i}d_{i}-b_{i}c_{i}=1$ for each $i\in\mathcal{I}_{N}$. Let $(i_{n})_{n}\in(\mathcal{I}_{N})^{\mathbb{N}}$. Let
@@
Let $C_{1}:=\max\{|a_{i}|,|b_{i}|,|c_{i}|,|d_{i}|:i\in\mathcal{I}_{N}\}$. Then, for every $n$, $|r_{n+1}|+|s_{n+1}|\leq C_{1}(|r_{n}|+|s_{n}|)$. Hence, $|r_{n}|+|s_{n}|\leq 2C_{1}^{n}$ for every $n$.
Assume that $\sigma=(i_{1},\cdots,i_{n})$. Then, $|\sigma|=n$ and
$h_{\sigma}(1)-h_{\sigma}(0)=\frac{1}{s_{n}(r_{n}+s_{n})}\geq\frac{1}{(|r_{n}|+|s_{n}|)^{2}}\geq\frac{1}{4|C_{1}|^{2n}}.$
@@
Let $\{h_{i}\}_{i}$ be an LF system. Then, there exists $C>0$ such that for every $\sigma\in(\mathcal{I}_{N})^{*}$,
@@
We first remark that
$(c_{i}x+d_{i})(c_{i}y+d_{i})\geq 1+|c_{i}||x-y|,\ x,y\in[0,1].$ (2.5)
We assume that $x\leq y$. Assume $c_{i}>0$. Then,
$c_{i}y+d_{i}\geq c_{i}x+d_{i}\geq\min_{x\in[0,1]}c_{i}x+d_{i}=\min\{d_{i},c_{i}+d_{i}\}=d_{i}\geq 1.$
@@
Assume $c_{i}<0$. Then,
$(c_{i}x+d_{i})(c_{i}y+d_{i})=(c_{i}y+d_{i})^{2}-c_{i}(y-x)(c_{i}x+d_{i})\geq 1-c_{i}(y-x).$
If $c_{i}=0$, then, $(c_{i}x+d_{i})(c_{i}y+d_{i})=d_{i}^{2}\geq 1=1+|c_{i}||x-y|$. Thus we see (2.5).
By (2.5),
@@
Since $h_{i}(1)-h_{i}(0)<1$, $a_{i}<d_{i}$. Let $C_{1}:=\min\{|c_{i}|:c_{i}\neq 0\}$ and $C_{2}:=\max\left\{\dfrac{a_{i}}{d_{i}}:c_{i}=0\right\}$. Then, $C_{1}>0$, $C_{2}<1$ and
$\max_{i}|h_{i}(x)-h_{i}(y)|\leq\max\left\{\frac{|x-y|}{1+C_{1}|x-y|},C_{2}|x-y|\right\}.$
@@
$\max_{i}|h_{i}(x)-h_{i}(y)|\leq\frac{|x-y|}{1+C_{3}|x-y|},\ x,y\in[0,1].$
Let $a_{n}:=\max\left\{h_{\sigma}(1)-h_{\sigma}(0):\sigma\in(\mathcal{I}_{N})^{*},|\sigma|=n\right\}$. Then, $a_{1}\leq\dfrac{1}{1+C_{3}}$ and $a_{n+1}\leq\dfrac{a_{n}}{1+C_{3}a_{n}},n\geq 1$. Hence,
$a_{n}\leq\frac{1}{nC_{3}+1}\leq\frac{1}{nC_{3}}.$
@@
Assume that $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ and $\{g_{i}\}_{i\in\mathcal{I}_{N}}$ are both LF systems and let $\varphi$ be the solution of (1.1). Then, there exists a constant $C>0$ such that for every $y\in(0,1)$ and $r>0$ satisfying that $0\leq y-r<y+r\leq 1$,
@@
This might be regarded as a weak version of the regularity of $\mu_{\varphi}$ in the sense of Ullman-Stahl-Totik. See [22, Criterion 1.1]. However, the following proof does not yield the regularity.
@@
We apply Lemma 2.10 to $\{g_{i}\}_{i}$ and Lemma 2.11 to $\{f_{i}\}_{i}$. Let $C$ be the constant appearing in Lemma 2.11. Take $y\in(0,1)$ and $r>0$ such that $0\leq y-r<y+r\leq 1$ arbitralily. Let $n$ be a natural number such that $\frac{C}{n}<r\leq\frac{C}{n-1}$. Then, by the compatibility condition and Lemma 2.11, there exist $i_{1},\cdots,i_{n}\in\mathcal{I}_{N}$ such that
$[f_{i_{1}}\circ\cdots\circ f_{i_{n}}(0),f_{i_{1}}\circ\cdots\circ f_{i_{n}}(1)]\subset[y-r,y+r].$
@@
$\mu_{\varphi}([y-r,y+r])\geq g_{i_{1}}\circ\cdots\circ g_{i_{n}}(1)-g_{i_{1}}\circ\cdots\circ g_{i_{n}}(0).$
Let $c$ be the constant appearing in Lemma 2.10. By Lemma 2.10,
$\mu_{\varphi}([y-r,y+r])\geq c^{n}\geq c^{1+C/r}.$

### v4

@@ 2.2. Weak contraction
Hereafter, we assume additional conditions for the D-systems $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ and $\{g_{i}\}_{i\in\mathcal{I}_{N}}$. We say that a function $h\colon[0,1]\to[0,1]$ is a contraction on $[0,1]$ if
$\|h\|_{\textup{Lip}}\coloneqq\sup_{x\neq y}\left|\frac{h(x)-h(y)}{x-y}\right|<1.$
$\|h\|_{\textup{Lip}}$ is called the Lipschitz constant of $h$.
For a function $h\colon[0,1]\to\mathbb{R}$, let
(2.5) $\omega_{h}(s)\coloneqq\sup\left\{|h(x)-h(y)|\colon|x-y|\leq s\right\}.$
We say that a function $h\colon[0,1]\to[0,1]$ is a weak contraction on $[0,1]$ if
$\phi_{h}(t)\coloneqq\lim_{s\to t+0}\omega_{h}(s)<t,\ \ t>0.$
@@
Let $h$ be a weak contraction. Then, $\displaystyle\lim_{n\to\infty}\phi_{h}^{n}(t)=0$ for every $t>0$ and $|h(x)-h(y)|\leq\phi_{h}(|x-y|)$ for $x,y\in[0,1]$.
@@
Since $\omega_{h}$ is increasing, $\phi_{h}$ is increasing and right-continuous on $[0,1]$. Hence,
$|h(x)-h(y)|\leq\lim_{s\to|x-y|+0}\phi_{h}(s)=\phi_{h}(|x-y|).$
By the assumption, for every $t>0$, $(\phi_{h}^{n}(t))_{n}$ is strictly decreasing. Hence, the limit $\displaystyle\phi^{\infty}_{h}(t)\coloneqq\lim_{n\to\infty}\phi_{h}^{n}(t)$ exists. By the right-continuity of $\phi_{h}$, we see that $\displaystyle\phi_{h}\left(\phi^{\infty}_{h}(t)\right)=\phi^{\infty}_{h}(t)$. Hence, $\phi^{\infty}_{h}(t)=0$. ∎
@@
Let $\{h_{i}\}_{i\in\mathcal{I}_{N}}$ be a compatible system on $[0,1]$ such that each $h_{i}$ is a weak contraction on $[0,1]$. Then, $\{h_{i}\}_{i\in\mathcal{I}_{N}}$ is a D-system.
@@
Let $\displaystyle\phi\coloneqq\max_{i\in\mathcal{I}_{N}}\phi_{h_{i}}$. Since each $\phi_{h_{i}}$ is right-continuous, $\phi$ is right-continuous [^1] The assumption that $N$ is finite is important. on $[0,1]$, $\phi(t)<t$ for every $t>0$, and $\displaystyle\max_{i\in\mathcal{I}_{N}}|h_{i}(x)-h_{i}(y)|\leq\phi(|x-y|)$, $x,y\in[0,1]$. Then, we see that $\displaystyle\lim_{n\to\infty}\phi^{n}(t)=0$ for every $t>0$. Hence,
@@
$\lim_{n\to\infty}\max_{i_{1},\ldots,i_{n}\in\mathcal{I}_{N}}\left|h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1)-h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0)\right|=0.$
@@
$0,1]=\bigcup_{i_{1},\ldots,i_{n}\in\mathcal{I}_{N}}\left[h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0),h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1)\right],$
@@
We say that a compatible system $\{h_{i}\}_{i\in\mathcal{I}_{N}}$ on $[0,1]$ is a (weakly) contractive system on $[0,1]$ if each $h_{i}$ is a (weak) contraction on $[0,1]$. Every weakly contractive system is a D-system.
@@ 2.3. Linear fractional transforms
We say that a family of functions $\{h_{i}\}_{i\in\mathcal{I}_{N}}$ is an LF system on $[0,1]$ if $h_{i}(x)\coloneqq\dfrac{a_{i}x+b_{i}}{c_{i}x+d_{i}},x\in[0,1]$, for each $i\in\mathcal{I}_{N}$, where $a_{i},b_{i},c_{i},d_{i},i\in\mathcal{I}_{N}$, are real numbers such that
(1) (well-defined) It holds that $c_{i}+d_{i}>0$ and $d_{i}>0$ for $i\in\mathcal{I}_{N}$.
(2) (strictly increasing, weak contraction) It holds that
@@
(3) (compatibility) It holds that $b_{0}=0$ and $\dfrac{a_{i}+b_{i}}{c_{i}+d_{i}}=\dfrac{b_{i+1}}{d_{i+1}}$ for $i\in\mathcal{I}_{N}$, where we let $\dfrac{b_{N}}{d_{N}}\coloneqq 1$.
(4) The sequence $\left(\dfrac{b_{i}}{d_{i}}\right)_{i}$ is strictly increasing in $i$.
Condition (4) follows from the conditions (1), (2) and (3) but is frequently used. For condition (2), we may consider a stronger condition:
@@
Since $h_{i}$ is invariant if we multiply each of $a_{i},b_{i},c_{i}$ and $d_{i}$ by a positive constant, we may assume that $d_{i}=1$ or $a_{i}d_{i}-b_{i}c_{i}=1$. An LF system is a compatible system.
@@
Every LF system is a weakly contractive system. If (2’) holds instead of (2), then it is a contractive system.
@@
Let $\{h_{i}\}_{i\in\mathcal{I}_{N}}$ be an LF-system. We assume that $d_{i}=1$ for every $i\in\mathcal{I}_{N}$. By Lemma 2.7, it suffices to show that each $h_{i}$ is a weak contraction. We first remark that
(2.6) $|h_{i}(x)-h_{i}(y)|=\frac{a_{i}-b_{i}c_{i}}{(c_{i}x+1)(c_{i}y+1)}|x-y|.$
@@
(iii) Assume $c_{i}=0$.
We remark that $0<a_{i}\leq 1$ and $|h_{i}(x)-h_{i}(y)|=a_{i}|x-y|$. By the assumption for $a_{i},b_{i},c_{i}$, it holds that $h_{j}(1)>h_{j}(0)$ for every $j$. By the fact that $N\geq 2$, we obtain that $a_{i}=h_{i}(1)-h_{i}(0)<1$. ∎
@@
Let $\displaystyle(\mathcal{I}_{N})^{*}\coloneqq\bigcup_{n\in\mathbb{N}}(\mathcal{I}_{N})^{n}$. The length of $\sigma$ is the unique number $n\in\mathbb{N}$ such that $\sigma\in(\mathcal{I}_{N})^{n}$. We denote the length of $\sigma\in(\mathcal{I}_{N})^{*}$ by $|\sigma|$. For an LF system $\{h_{i}\}_{i\in\mathcal{I}_{N}}$ and $\sigma=(\sigma_{1},\ldots,\sigma_{n})\in(\mathcal{I}_{N})^{*}$, we let
$h_{\sigma}(x)\coloneqq h_{\sigma_{1}}\circ\cdots\circ h_{\sigma_{n}}(x),\ x\in[0,1].$
@@
Let $\{h_{i}\}_{i\in\mathcal{I}_{N}}$ be an LF system. Then, there exists $c\in(0,1)$ such that for every $\sigma\in(\mathcal{I}_{N})^{*}$,
@@
Let $A_{i}\coloneqq\begin{pmatrix}a_{i}&b_{i}\\ c_{i}&d_{i}\end{pmatrix}$, $i\in\mathcal{I}_{N}$. We assume that $a_{i}d_{i}-b_{i}c_{i}=1$ for each $i\in\mathcal{I}_{N}$. Let $(i_{n})_{n}\in(\mathcal{I}_{N})^{\mathbb{N}}$ and
@@
Let $C_{1}\coloneqq\max\{|a_{i}|,|b_{i}|,|c_{i}|,|d_{i}|\colon i\in\mathcal{I}_{N}\}$. Then, for every $n$, $|r_{n+1}|+|s_{n+1}|\leq 2C_{1}(|r_{n}|+|s_{n}|)$. Hence, $|r_{n}|+|s_{n}|\leq(2C_{1})^{n}$ for every $n$.
Assume that $\sigma=(i_{1},\ldots,i_{n})$. Then, $|\sigma|=n$ and by using the equality $p_{n}s_{n}-q_{n}r_{n}=1$,
$h_{\sigma}(1)-h_{\sigma}(0)=\frac{p_{n}+q_{n}}{r_{n}+s_{n}}-\frac{q_{n}}{s_{n}}$
$\frac{1}{s_{n}(r_{n}+s_{n})}\geq\frac{1}{(|r_{n}|+|s_{n}|)^{2}}\geq\frac{1}{4^{n}|C_{1}|^{2n}}.$
@@
Let $\{h_{i}\}_{i\in\mathcal{I}_{N}}$ be an LF system. Then, there exists $C>0$ such that for every $\sigma\in(\mathcal{I}_{N})^{*}$,
@@
In this proof, we assume that $a_{i}d_{i}-b_{i}c_{i}=1$ for each $i\in\mathcal{I}_{N}$. We first show that
(2.7) $(c_{i}x+d_{i})(c_{i}y+d_{i})\geq 1+|c_{i}||x-y|,\ x,y\in[0,1].$
We assume that $x\leq y$. Assume $c_{i}>0$. Then, by Definition 2.8 (2),
$c_{i}y+d_{i}\geq c_{i}x+d_{i}\geq\min_{z\in[0,1]}c_{i}z+d_{i}=d_{i}\geq 1.$
@@
Assume $c_{i}<0$. Then, by Definition 2.8 (2) and (1),
$c_{i}x+d_{i}\geq c_{i}y+d_{i}\geq\min_{z\in[0,1]}c_{i}z+d_{i}=c_{i}+d_{i}=|c_{i}+d_{i}|\geq 1.$
Hence,
$(c_{i}x+d_{i})(c_{i}y+d_{i})=(c_{i}y+d_{i})^{2}-c_{i}(y-x)(c_{i}y+d_{i})\geq 1-c_{i}(y-x).$
If $c_{i}=0$, then $(c_{i}x+d_{i})(c_{i}y+d_{i})=d_{i}^{2}\geq 1=1+|c_{i}||x-y|$. Thus we see (2.7).
By (2.6) and (2.7),
@@
Since $h_{i}(1)-h_{i}(0)<1$, it holds that $a_{i}<d_{i}$. Let $C_{1}\coloneqq\min\{|c_{i}|\colon c_{i}\neq 0\}$ and $C_{2}\coloneqq\max\left\{\dfrac{a_{i}}{d_{i}}\colon c_{i}=0\right\}$. Then, $C_{1}>0$, $C_{2}<1$ and
$\max_{i\in\mathcal{I}_{N}}|h_{i}(x)-h_{i}(y)|\leq\max\left\{\frac{|x-y|}{1+C_{1}|x-y|},C_{2}|x-y|\right\}.$
@@
$\max_{i\in\mathcal{I}_{N}}|h_{i}(x)-h_{i}(y)|\leq\frac{|x-y|}{1+C_{3}|x-y|},\ x,y\in[0,1].$
Let $A_{n}\coloneqq\max\left\{h_{\sigma}(1)-h_{\sigma}(0)\colon\sigma\in(\mathcal{I}_{N})^{*},|\sigma|=n\right\}$. Then, $A_{1}\leq\dfrac{1}{1+C_{3}}$ and $A_{n+1}\leq\dfrac{A_{n}}{1+C_{3}A_{n}},n\geq 1$. Hence,
$A_{n}\leq\frac{1}{nC_{3}+1}\leq\frac{1}{nC_{3}}.\qed$
@@
Assume that $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ and $\{g_{i}\}_{i\in\mathcal{I}_{N}}$ are LF systems and let $\varphi$ be the solution of (1.1). Then, there exists a constant $C>0$ such that for every $y\in(0,1)$ and $r>0$ satisfying that $0\leq y-r<y+r\leq 1$,
@@
This might be regarded as a weak version of the regularity of $\mu_{\varphi}$ in the sense of Ullman–Stahl–Totik. See [22, Criterion 1.1]. However, the following proof does not yield the regularity.
@@
We apply Lemma 2.11 to $\{g_{i}\}_{i\in\mathcal{I}_{N}}$ and Lemma 2.12 to $\{f_{i}\}_{i\in\mathcal{I}_{N}}$. Let $C$ be the constant appearing in Lemma 2.12. Take $y\in(0,1)$ and $r>0$ such that $0\leq y-r<y+r\leq 1$ arbitrarily. Let $n$ be a natural number such that $\dfrac{C}{n}<r\leq\dfrac{C}{n-1}$. We let $\dfrac{C}{n-1}\coloneqq\infty$ for $n=1$. Then, by the compatibility condition and Lemma 2.12, there exist $i_{1},\ldots,i_{n}\in\mathcal{I}_{N}$ such that
$\left[f_{i_{1}}\circ\cdots\circ f_{i_{n}}(0),f_{i_{1}}\circ\cdots\circ f_{i_{n}}(1)\right]\subset[y-r,y+r].$
@@
$\mu_{\varphi}\left([y-r,y+r]\right)\geq\mu_{\varphi}\left(\left[f_{i_{1}}\circ\cdots\circ f_{i_{n}}(0),f_{i_{1}}\circ\cdots\circ f_{i_{n}}(1)\right]\right)$
$\varphi(f_{i_{1}}\circ\cdots\circ f_{i_{n}}(1))-\varphi(f_{i_{1}}\circ\cdots\circ f_{i_{n}}(0))$
$g_{i_{1}}\circ\cdots\circ g_{i_{n}}(1)-g_{i_{1}}\circ\cdots\circ g_{i_{n}}(0).$
By Lemma 2.11, there exists a constant $c\in(0,1)$ depending only on $\{g_{i}\}_{i}$ such that
$g_{i_{1}}\circ\cdots\circ g_{i_{n}}(1)-g_{i_{1}}\circ\cdots\circ g_{i_{n}}(0)\geq c^{n}.$
By the assumption for $n$, we obtain that $c^{n}\geq c^{1+C/r}$. Thus, we obtain that
$\mu_{\varphi}([y-r,y+r])\geq c^{1+C/r}.\qed$